# Fairness-Regulated Dense Subgraph Discovery

### v1

@@ 7.2 Evaluation of FADSG-II
Direct comparison with prior art is not feasible here, since FADSG-II utilizes a different fairness criterion, as detailed in the “Problem Statement” section. In Figure 7, we compare this formulation with DSG. The top section depicts the PoF of the solution obtained from FADSG-II, and the bottom one illustrates the fraction of all of the protected vertices in the solutions of DSG and FADSG-II. For the Twitch datasets, DSG’s solution includes only about $1/5$ of the protected vertices. This is clearly due to the fact that DSG does not take into account the protected attributes of the vertices. On the other hand, FADSG-II for $\delta=1$, yields solutions with nearly half of the protected vertices, while achieving a relatively low PoF. Similar to FADSG-I, there are instances where the solution does not precisely align with the desired fairness level. This can be attributed to the structure of each graph, which may not inherently exhibit the specific fraction we are aiming for (see Appendix I for more details).
@@ 8 Conclusions
We addressed the fairness-aware densest subgraph problem by introducing two tractable formulations, each incorporating a different regularizer in the objective function, to provide two different notions of fairness. The inclusion of fairness introduces the inherent fairness-density trade-off. In order to quantify this trade-off we proposed a measure of the price of fairness, and we analyzed it based on an illuminating example – which revealed the underlying tension between density and fairness. To assess the effectiveness of the proposed solutions, we employed a peeling-based approximation algorithm, and compared their performance against the prior art using real-world datasets. Our findings revealed that FADSG-I generally incurs a lower price of fairness (PoF) in achieving perfect balanced subgraphs, sometimes even less than half of that of prior art. FADSG-I can also induce a perfectly balanced subgraph in some extreme scenarios where existing approaches fall short, and it is capable of providing various fairness levels, other than perfect balance. Furthermore, FADSG-II promotes an alternative notion of inclusion of protected vertices, often achieving this with a relatively low PoF. Together, these formulations offer flexible solutions for balancing fairness and density in subgraph discovery.
@@ A.1 Proof of Lemma 1
The claim then follows by noting that
$r_{2}(\mathcal{S})=1\Leftrightarrow|\mathcal{S}_{p}|-2|\mathcal{S}\cap\mathcal{S}_{p}|=0.$ (9)
A.2 Proof of Lemma 2
@@
$r_{2}(\mathcal{S})=1+\frac{|\mathcal{S}_{p}|}{|\mathcal{S}|}-2r_{1}(\mathcal{S}).$ (10)
Since $|\mathcal{S}_{p}|\geq|\mathcal{S}_{p}\cap\mathcal{S}|=r_{1}(\mathcal{S})|\mathcal{S}|$, it implies that $\frac{|\mathcal{S}_{p}|}{|\mathcal{S}|}\geq r_{1}(\mathcal{S})$. Hence, we obtain
@@ Appendix B Tractability details
First, consider problem (1). Define the functions $p_{1}({\mathbf{x}},\lambda):={\mathbf{x}}^{\top}{\mathbf{A}}{\mathbf{x}}+\lambda\cdot{\mathbf{e}}_{p}^{\top}{\mathbf{x}}$, and $q({\mathbf{x}}):={\mathbf{e}}^{\top}{\mathbf{x}}$. Thus, the objective function can be expressed as $g({\mathbf{x}},\lambda)=p_{1}({\mathbf{x}},\lambda)/q({\mathbf{x}})$. Since $p_{1}(\cdot,\lambda)$ is the sum of a non-negative, monotone supermodular function and a non-negative modular function, the result is also non-negative, monotone supermodular (for every $\lambda\geq 0$). This implies that FADSG-I corresponds to maximizing the ratio of a monotone, non-negative supermodular function and a modular function. The solution to such a problem can be determined in polynomial-time by solving a finite number of supermodular maximization (SFM) sub-problems within a binary-search framework [48, 49]. Since SFM incurs polynomial-time complexity [50], the claim follows. Furthermore, the quadratic form of $p_{1}(\cdot,\lambda)$ can be exploited to show that each sub-problem is equivalent to solving a minimum-cut problem [14, 51], which can be accomplished via maximum-flow. Each maximum-flow problem incurs complexity $O(nm)$ [52] and an upper-bound on the maximum number of binary-search steps required is $O(\log(n+\lambda)+\log(n))$ (see Appendix B.1).
@@
Next, consider problem (3), and define the function $p_{2}({\mathbf{x}},\lambda):={\mathbf{x}}^{\top}{\mathbf{A}}{\mathbf{x}}-\lambda\cdot({\mathbf{e}}-2{\mathbf{e}}_{p})^{\top}{\mathbf{x}}-\lambda\cdot n_{p}$. Then, the objective function of FADSG-II can be expressed as the ratio $h({\mathbf{x}},\lambda)=p_{2}({\mathbf{x}},\lambda)/q({\mathbf{x}})$. While $p_{2}(\cdot,\lambda)$ is a supermodular function (being the sum of a supermodular and a modular function), it is not guaranteed to be non-negative for every choice of $\lambda>0$. Additionally, it is not normalized either, since $p_{2}(\mathbf{0},\lambda)=-\lambda\cdot n_{p}$. Nevertheless, we can still show that (3) can be solved via a sequence of maximum-flow problems using a small modification of the binary-search strategy explained before for FADSG-I (see Appendix C for full details). Thus, FADSG-II can also be solved exactly in polynomial-time.
B.1 Upper bound on the number of binary search iterations for FADSG-I and FADSG-II

### v2

@@ 7.2 Evaluation of FADSG-II
Direct comparison with prior art is not feasible here, since FADSG-II utilizes a different fairness criterion, as detailed in Section 3. In Figure 8, we compare this formulation with DSG. The top section depicts the PoF of the solution obtained from FADSG-II, and the bottom one illustrates the fraction of all of the protected vertices in the solutions of DSG and FADSG-II. For the Twitch datasets, DSG’s solution includes only about $1/5$ of the protected vertices. This is clearly due to the fact that DSG does not take into account the protected attributes of the vertices. On the other hand, FADSG-II for $\delta=1$, yields solutions with nearly half of the protected vertices, while achieving a relatively low PoF. Similar to FADSG-I, there are instances where the solution does not precisely align with the desired fairness level. This can be attributed to the structure of each graph, which may not inherently exhibit the specific fraction we are aiming for (see Appendix H for more details).
@@ 8 Conclusions
We addressed the fairness-aware densest subgraph problem by introducing two tractable formulations, each incorporating a different regularizer in the objective function, to provide two different notions of fairness. The inclusion of fairness introduces the inherent fairness-density trade-off. In order to quantify this trade-off we proposed a measure of the price of fairness, and we analyzed it based on an illuminating example – which revealed the underlying tension between density and fairness. To assess the effectiveness of the proposed solutions, we employed a peeling-based approximation algorithm, and compared their performance against the prior art using real-world datasets. Our findings revealed that FADSG-I generally incurs a lower price of fairness (PoF) in achieving perfect balanced subgraphs, sometimes even less than half of that of prior art. FADSG-I can also induce a perfectly balanced subgraph in some extreme scenarios where existing approaches fall short, and it is capable of providing various fairness levels, other than perfect balance. Furthermore, FADSG-II promotes an alternative notion of inclusion of protected vertices, often achieving this with a relatively low PoF. Together, these formulations offer flexible solutions for balancing fairness and density in subgraph discovery. Extending our approach to handle multiple vertex groups using regularizers which are designed to ensure that no protected group is inadequately represented constitutes a direction of future work.
Broader Impact
The authors posit this work on the assumption that practitioners will appropriately select the protected subset and publicly disclose it along with the results of their analysis.
Acknowledgments
Supported in part by the National Science Foundation, USA under grant IIS-1908070 and the KU Leuven Special Research Fund BOF/STG-22-040.
@@ A.1 Proof of Lemma 4
The claim then follows by noting that $r_{2}(\mathcal{S})=1\Leftrightarrow|\mathcal{S}_{p}|-2|\mathcal{S}\cap\mathcal{S}_{p}|=0$. ∎
A.2 Proof and Intuition of Lemma 5
@@
$r_{2}(\mathcal{S})=1+\frac{n_{p}}{|\mathcal{S}|}-2r_{1}(\mathcal{S}).$ (10)
$\bullet$ For the first inequality: Since $|\mathcal{S}_{p}|\geq|\mathcal{S}_{p}\cap\mathcal{S}|=r_{1}(\mathcal{S})|\mathcal{S}|$, it implies that $\frac{|\mathcal{S}_{p}|}{|\mathcal{S}|}\geq r_{1}(\mathcal{S})$. Hence, we obtain
@@
$\bullet$ For the second inequality: We have that $\frac{n_{p}}{|\mathcal{S}|}\leq\frac{n_{p}}{2}$, since each solution has at least one edge. Hence, we obtain
$2r_{1}(\mathcal{S})+r_{2}(\mathcal{S})\leq 1+\frac{n_{p}}{2}.$ (12)
Figure 9: The joint fairness region of the regularizers $r_{1}(\mathcal{S})$ and $r_{2}(\mathcal{S})$, for a graph with $n_{p}$ protected vertices.
Furthermore, we define the joint fairness region characterized by the above inequalities as the following polytope:
:=$\!\!\!\mathcal{R}\!:=\!\left\{(r_{1}(\mathcal{S}),r_{2}(\mathcal{S})):0\leq r_{1}(\mathcal{S})\leq 1,0\leq r_{2}(\mathcal{S})\leq 1+\frac{n_{p}}{2},r_{1}(\mathcal{S})+r_{2}(\mathcal{S})\geq 1,2r_{1}(\mathcal{S})+r_{2}(\mathcal{S})\leq 1+\frac{n_{p}}{2}\right\}$ (13)
$\setminus\{r_{1}(\mathcal{S})=0\land r_{2}(\mathcal{S})=1\}$
Note that the point $(0,1)$ is excluded from $\mathcal{R}$, since if $r_{1}(\mathcal{S})=0\Leftrightarrow|\mathcal{S}\cap\mathcal{S}_{p}|=0$ and $r_{2}(\mathcal{S})=1\Leftrightarrow|\mathcal{S}\cap\mathcal{S}_{p}|=\frac{1}{2}|\mathcal{S}_{p}|\Rightarrow|\mathcal{S}_{p}|=0$, which is a contradiction since we assume that there always exists at least one protected vertex in the graph. We can approximate this point when $n_{p}\ll n$. In this case, when $r_{1}(\mathcal{S})\rightarrow 0$, it can still hold that half of the protected vertices belong in the solution, i.e., $r_{2}(\mathcal{S})=1$.
Let us consider the three remaining corner-points of $\mathcal{R}$.
  • (Bottom right:) We can attain $(1,0)$ when $r_{2}(\mathcal{S})=0$, meaning that the solution is the protected set $\mathcal{S}_{p}$, hence $r_{1}(\mathcal{S})=\frac{|\mathcal{S}\cap\mathcal{S}_{p}|}{|\mathcal{S}|}=1$.
  • (Top left:) We can attain $(0,\frac{n_{p}+2}{2})$ only when the solution contains just one edge that connects two unprotected vertices. This holds because when $r_{1}(\mathcal{S})=0\Leftrightarrow|\mathcal{S}\cap\mathcal{S}_{p}|=0\Rightarrow r_{2}(\mathcal{S})=\frac{n_{p}}{2}+1$, which is equal to the desired value only for $|\mathcal{S}|=2$.
  • (Top right:) We can attain $(1,\frac{n_{p}-2}{2})$ only when the solution contains just one edge that connects two protected vertices. This holds because when $r_{1}(\mathcal{S})=1\Leftrightarrow|\mathcal{S}\cap\mathcal{S}_{p}|=|\mathcal{S}|\Rightarrow r_{2}(\mathcal{S})=\frac{n_{p}}{2}-1$, which is equal to the desired value only for $|\mathcal{S}|=2\Leftrightarrow|\mathcal{S}\cap\mathcal{S}_{p}|=2$.
@@ Appendix B Tractability details
First, consider problem (2). Define the functions $p_{1}({\mathbf{x}},\lambda):={\mathbf{x}}^{\top}{\mathbf{A}}{\mathbf{x}}+\lambda\cdot{\mathbf{e}}_{p}^{\top}{\mathbf{x}}$, and $q({\mathbf{x}}):={\mathbf{e}}^{\top}{\mathbf{x}}$. Thus, the objective function can be expressed as $g({\mathbf{x}},\lambda)=p_{1}({\mathbf{x}},\lambda)/q({\mathbf{x}})$. Since $p_{1}(\cdot,\lambda)$ is the sum of a non-negative, monotone supermodular function and a non-negative modular function, the result is also non-negative, monotone supermodular (for every $\lambda\geq 0$). This implies that FADSG-I corresponds to maximizing the ratio of a monotone, non-negative supermodular function and a modular function. The solution to such a problem can be determined in polynomial-time by solving a finite number of supermodular maximization (SFM) sub-problems within a binary-search framework (Fujishige, 2005; Bai et al., 2016). Since SFM incurs polynomial-time complexity (Grötschel et al., 2012), the claim follows. Furthermore, the quadratic form of $p_{1}(\cdot,\lambda)$ can be exploited to show that each sub-problem is equivalent to solving a minimum-cut problem (Goldberg, 1984; Kolmogorov & Zabih, 2002), which can be accomplished via maximum-flow, as we show in Appendix B.1.1. Each maximum-flow problem incurs complexity $O(nm)$ (Orlin, 2013) and an upper-bound on the maximum number of binary-search steps required is $O(\log(n+\lambda)+\log(n))$ (see Appendix B.2).
@@
Next, consider problem (4), and define the function $p_{2}({\mathbf{x}},\lambda):={\mathbf{x}}^{\top}{\mathbf{A}}{\mathbf{x}}-\lambda\cdot({\mathbf{e}}-2{\mathbf{e}}_{p})^{\top}{\mathbf{x}}-\lambda\cdot n_{p}$. Then, the objective function of FADSG-II can be expressed as the ratio $h({\mathbf{x}},\lambda)=p_{2}({\mathbf{x}},\lambda)/q({\mathbf{x}})$. While $p_{2}(\cdot,\lambda)$ is a supermodular function (being the sum of a supermodular and a modular function), it is not guaranteed to be non-negative for every choice of $\lambda>0$. Additionally, it is not normalized either, since $p_{2}(\mathbf{0},\lambda)=-\lambda\cdot n_{p}$. Nevertheless, we can still show that (4) can be solved via a sequence of maximum-flow problems using a small modification of the binary-search strategy, as we explain in Appendix B.1.2. Thus, FADSG-II can also be solved exactly in polynomial-time.
B.1 Exact solutions via max-flow
B.1.1 Applying max-flow on FADSG-I
Figure 10: An edge-weighted directed graph $(\mathcal{U},\mathcal{A},w_{\beta})$ constructed from $\mathcal{G}$, $\beta$ and $\lambda$.
Based on the interpretation of Lanciano et al. (2024, Algorithm 1) for Goldberg’s max-flow based algorithm (Goldberg, 1984), we only need to make a small modification in order to apply it to FADSG-I.
The algorithm maintains an estimate of the unknown optimal value by tracking upper and lower bounds, which it progressively tightens through binary search. To update them, the algorithm employs a max-flow (equivalently min-cut) computation. Let $\beta\geq 0$ be the midpoint of the bounds kept in the current iteration of the max-flow based algorithm. For $\mathcal{G}=(\mathcal{V},\mathcal{E})$, $\mathcal{S}_{p}$ the protected set and $\lambda$ the regularization parameter of FADSG-I (constant), the algorithm constructs the following edge-weighted directed graph $(\mathcal{U},\mathcal{A},w^{I}_{\beta})$ (Figure 10), with $\mathcal{U}=\mathcal{V}\cup\{s,t\},\ \mathcal{A}=\mathcal{A}_{s}\cup\mathcal{A}_{\mathcal{E}}\cup\mathcal{A}^{p}_{t}\cup\mathcal{A}^{\bar{p}}_{t}$, where $\mathcal{A}_{s}=\{(s,v)|v\in\mathcal{V}\}$, $\mathcal{A}_{\mathcal{E}}=\{(u,v),(v,u)|\{u,v\}\in\mathcal{E}\}$, $\mathcal{A}_{t}^{p}=\{(u,t)|v\in\mathcal{S}_{p}\}$ and $\mathcal{A}_{t}^{\bar{p}}=\{(u,t)|v\in\bar{\mathcal{S}_{p}}\}$ and $w^{I}_{\beta}:\mathcal{A}\rightarrow\mathbb{R}_{+}$ such that
$w^{I}_{\beta}(e)=\left\{\begin{array}[]{ll}\text{deg}(v)&e=(s,v)\in\mathcal{A}_{s}\\ 1&e\in\mathcal{A}_{\mathcal{E}}\\ \beta-\lambda&e\in\mathcal{A}_{t}^{p}\\ \beta&e\in\mathcal{A}_{t}^{\bar{p}}\\ \end{array}\right..$ (14)
An $s-t$ cut of $(\mathcal{U},\mathcal{A},w_{\beta})$ is a partition of $(\mathcal{X},\mathcal{Y})$ of $\mathcal{U}$, such that $s\in\mathcal{X}$ and $t\in\mathcal{Y}$. Then define the cost of an $s-t$ cut $(\mathcal{X},\mathcal{Y})$, as ${\mathrm{cost}(\mathcal{X},\mathcal{Y}):=\sum_{(u,v)\in\mathcal{A}:u\in\mathcal{X},v\in\mathcal{Y}}w_{\beta}(u,v)}$.
Lemma 8.
Let $(\mathcal{X},\mathcal{Y})$ be an $s-t$ cut of the edge-weighted directed graph $(\mathcal{U},\mathcal{A},w^{I}_{\beta})$, and $\mathcal{S}=\mathcal{X}\setminus\{s\}$. Then it holds that $\mathrm{cost}(\mathcal{X},\mathcal{Y})=2m+\beta|\mathcal{S}|-p_{1}(\mathcal{S},\lambda)$, where $p_{1}(\mathcal{S},\lambda)=2e(\mathcal{S})+\lambda|\mathcal{S}\cap\mathcal{S}_{p}|$.
Proof.
Any edge from $\mathcal{X}$ to $\mathcal{Y}$ is contained in exactly one of the following sets: $\{(s,v)\in\mathcal{A}_{s}|v\in\mathcal{V}\setminus\mathcal{S}\}$, $\{(u,v)\in\mathcal{A}_{\mathcal{E}}|u\in\mathcal{S},v\in\mathcal{V}\setminus\mathcal{S}\}$, $\{(v,t)\in\mathcal{A}_{t}^{p}|v\in\mathcal{S}\}$, $\{(v,t)\in\mathcal{A}_{t}^{\bar{p}}|v\in\mathcal{S}\}$. Therefore the cost of $(\mathcal{X},\mathcal{Y})$ is
$\mathrm{cost}(\mathcal{X},\mathcal{Y})$ $=\sum_{(s,v)\in\mathcal{A}_{s}:v\in\mathcal{V}\setminus\mathcal{S}}\!\!\!\!\!\!\!\!w^{I}_{\beta}(s,v)+\!\!\!\!\sum_{(u,v)\in\mathcal{A}_{\mathcal{E}}:u\in\mathcal{S},v\in\mathcal{V}\setminus\mathcal{S}}\!\!\!\!\!\!\!\!w^{I}_{\beta}(u,v)+\!\!\!\!\sum_{(u,t)\in\mathcal{A}_{t}^{p}:v\in\mathcal{S}}\!\!\!\!\!\!w^{I}_{\beta}(v,t)+\!\!\!\!\sum_{(u,t)\in\mathcal{A}_{t}^{\bar{p}}:v\in\mathcal{S}}\!\!\!\!\!\!w^{I}_{\beta}(v,t)$ (15)
$=\sum_{v\in\mathcal{V}\setminus\mathcal{S}}\mathrm{deg}(v)+|\{(u,v)\in\mathcal{E}|u\in\mathcal{S},v\in\mathcal{V}\setminus\mathcal{S}\}+(\beta-\lambda)|\mathcal{S}\cap\mathcal{S}_{p}|+\beta(|\mathcal{S}|-|\mathcal{S}\cap\mathcal{S}_{p}|)$
$=2m-2e(\mathcal{S})+\beta|\mathcal{S}|-\lambda|\mathcal{S}\cap\mathcal{S}_{p}|=2m+\beta|\mathcal{S}|-2e(\mathcal{S})-\lambda|\mathcal{S}\cap\mathcal{S}_{p}|$
$=2m+\beta|\mathcal{S}|-p_{1}(\mathcal{S},\lambda).$
Let $(\mathcal{X},\mathcal{Y})$ be the minimum $s-t$ cut (or max-flow) of $(\mathcal{U},\mathcal{A},w^{I}_{\beta})$. If $\mathcal{X}\neq\{s\}$, then $\mathcal{S}=\mathcal{X}\setminus\{s\}$ is a vertex subset that satisfies $\beta|\mathcal{S}|-p_{1}(\mathcal{S},\lambda)\leq 0\Leftrightarrow\frac{p_{1}(\mathcal{S},\lambda)}{|\mathcal{S}|}\geq\beta$, and we can update the lower bound by $\beta$. If $\mathcal{X}=\{s\}$, then there is no $\mathcal{S}\subset\mathcal{V}$ that satisfies $\frac{p_{1}(\mathcal{S},\lambda)}{|\mathcal{S}|}>\beta$, and the upper bound can be replaced by $\beta$. Therefore, we can apply the same procedure as that of Lanciano et al. (2024, Algorithm 1), with a small change at the initial upper bound $\beta_{\mathrm{ub}}^{(0)}\leftarrow m$.
B.1.2 Applying max-flow on FADSG-II
We follow the same procedure as we did in Appendix B.1.1, but with a slightly different edge-weighted graph. For $\mathcal{G}=(\mathcal{V},\mathcal{E})$, $\mathcal{S}_{p}$ the protected set and $\lambda$ the regularization parameter of FADSG-II (constant), the algorithm constructs the following edge-weighted directed graph $(\mathcal{U},\mathcal{A},w^{II}_{\beta})$, with $\mathcal{U}=\mathcal{V}\cup\{s,t\},\ \mathcal{A}=\mathcal{A}_{s}\cup\mathcal{A}_{\mathcal{E}}\cup\mathcal{A}^{p}_{t}\cup\mathcal{A}^{\bar{p}}_{t}$, where $\mathcal{A}_{s}=\{(s,v)|v\in\mathcal{V}\}$, $\mathcal{A}_{\mathcal{E}}=\{(u,v),(v,u)|\{u,v\}\in\mathcal{E}\}$, $\mathcal{A}_{t}^{p}=\{(u,t)|v\in\mathcal{S}_{p}\}$ and $\mathcal{A}_{t}^{\bar{p}}=\{(u,t)|v\in\bar{\mathcal{S}_{p}}\}$ and $w^{II}_{\beta}:\mathcal{A}\rightarrow\mathbb{R}_{+}$ such that
$w^{II}_{\beta}(e)=\left\{\begin{array}[]{ll}\text{deg}(v)&e=(s,v)\in\mathcal{A}_{s}\\ 1&e\in\mathcal{A}_{\mathcal{E}}\\ \beta-\lambda&e\in\mathcal{A}_{t}^{p}\\ \beta+\lambda&e\in\mathcal{A}_{t}^{\bar{p}}\\ \end{array}\right..$ (16)
Lemma 9.
Let $(\mathcal{X},\mathcal{Y})$ be an $s-t$ cut of the edge-weighted directed graph $(\mathcal{U},\mathcal{A},w^{II}_{\beta})$, and $\mathcal{S}=\mathcal{X}\setminus\{s\}$. Then it holds that $\mathrm{cost}(\mathcal{X},\mathcal{Y})=(2m-\lambda\cdot n_{p})+\beta|\mathcal{S}|-p_{2}(\mathcal{S},\lambda)$, where $p_{2}(\mathcal{S},\lambda)=2e(\mathcal{S})-\lambda(|\mathcal{S}|+|\mathcal{S}_{p}|-2|\mathcal{S}\cap\mathcal{S}_{p}|)$.
Proof.
Any edge from $\mathcal{X}$ to $\mathcal{Y}$ is contained in exactly one of the following sets: $\{(s,v)\in\mathcal{A}_{s}|v\in\mathcal{V}\setminus\mathcal{S}\}$, $\{(u,v)\in\mathcal{A}_{\mathcal{E}}|u\in\mathcal{S},v\in\mathcal{V}\setminus\mathcal{S}\}$, $\{(v,t)\in\mathcal{A}_{t}^{p}|v\in\mathcal{S}\}$, $\{(v,t)\in\mathcal{A}_{t}^{\bar{p}}|v\in\mathcal{S}\}$. Therefore the cost of $(\mathcal{X},\mathcal{Y})$ is
$\mathrm{cost}(\mathcal{X},\mathcal{Y})$ $=\sum_{(s,v)\in\mathcal{A}_{s}:v\in\mathcal{V}\setminus\mathcal{S}}\!\!\!\!\!\!\!\!w^{II}_{\beta}(s,v)+\!\!\!\!\sum_{(u,v)\in\mathcal{A}_{\mathcal{E}}:u\in\mathcal{S},v\in\mathcal{V}\setminus\mathcal{S}}\!\!\!\!\!\!\!\!w^{II}_{\beta}(u,v)+\!\!\!\!\sum_{(u,t)\in\mathcal{A}_{t}^{p}:v\in\mathcal{S}}\!\!\!\!\!\!w^{II}_{\beta}(v,t)+\!\!\!\!\sum_{(u,t)\in\mathcal{A}_{t}^{\bar{p}}:v\in\mathcal{S}}\!\!\!\!\!\!w^{II}_{\beta}(v,t)$ (17)
$=\!\sum_{v\in\mathcal{V}\setminus\mathcal{S}}\!\!\mathrm{deg}(v)+|\{(u,v)\in\mathcal{E}|u\in\mathcal{S},v\in\mathcal{V}\setminus\mathcal{S}\}+(\beta-\lambda)|\mathcal{S}\cap\mathcal{S}_{p}|+(\beta+\lambda)(|\mathcal{S}|-|\mathcal{S}\cap\mathcal{S}_{p}|)$
$=2m+\beta|\mathcal{S}|-2e(\mathcal{S})+\lambda|\mathcal{S}|-2\lambda|\mathcal{S}\cap\mathcal{S}_{p}|=2m+\beta|\mathcal{S}|-(p_{2}(\mathcal{S},\lambda)+\lambda|\mathcal{S}_{p}|)$
$=(2m-\lambda\cdot n_{p})+\beta|\mathcal{S}|-p_{2}(\mathcal{S},\lambda).$
Similarly to FADSG-I, let $(\mathcal{X},\mathcal{Y})$ be the minimum $s-t$ cut (or max-flow) of $(\mathcal{U},\mathcal{A},w^{II}_{\beta})$. If $\mathcal{X}\neq\{s\}$, then $\mathcal{S}=\mathcal{X}\setminus\{s\}$ is a vertex subset that satisfies $\beta|\mathcal{S}|-p_{2}(\mathcal{S},\lambda)\leq 0\Leftrightarrow\frac{p_{2}(\mathcal{S},\lambda)}{|\mathcal{S}|}\geq\beta$, and we can update the lower bound by $\beta$. If $\mathcal{X}=\{s\}$, then there is no $\mathcal{S}\subset\mathcal{V}$ that satisfies $\frac{p_{2}(\mathcal{S},\lambda)}{|\mathcal{S}|}>\beta$, and the upper bound can be replaced by $\beta$. Therefore, we can again apply the same procedure as that of Lanciano et al. (2024, Algorithm 1), with a small change in the initial upper bound $\beta_{\mathrm{ub}}^{(0)}\leftarrow(m-\lambda\cdot n_{p})/2$, since $\lambda\cdot n_{p}$ is a constant.
B.2 Upper bound on the number of binary search iterations for FADSG-I and II